# Edge Delayed Deep Deterministic Policy Gradient: efficient continuous control for edge scenarios

### v1

@@ Abstract
Deep Reinforcement Learning is gaining increasing attention thanks to its capability to learn complex policies in high-dimensional settings. Recent advancements utilize a dual-network architecture to learn optimal policies through the Q-learning algorithm. However, this approach has notable drawbacks, such as an overestimation bias that can disrupt the learning process and degrade the performance of the resulting policy. To address this, novel algorithms have been developed that mitigate overestimation bias by employing multiple Q-functions.
Edge scenarios, which prioritize privacy, have recently gained prominence. In these settings, limited computational resources pose a significant challenge for complex Machine Learning approaches, making the efficiency of algorithms crucial for their performance.
In this work, we introduce a novel Reinforcement Learning algorithm tailored for edge scenarios, called Edge Delayed Deep Deterministic Policy Gradient (EdgeD3). EdgeD3 enhances the Deep Deterministic Policy Gradient (DDPG) algorithm, achieving significantly improved performance with $25\%$ less Graphics Process Unit (GPU) time while maintaining the same memory usage. Additionally, EdgeD3 consistently matches or surpasses the performance of state-of-the-art methods across various benchmarks, all while using $30\%$ fewer computational resources and requiring $30\%$ less memory.
keywords:
Edge Delayed Deep Deterministic Policy Gradient , Deep Reinforcement learning , Edge Computing , Q-Learning
journal: Engineering Applications of Artificial Intelligence
[…]
[inst1]organization=Human-Inspired Technology Research Center, University of Padova,addressline=Via Luzzatti, 4, city=Padova, postcode=35121, country=Italy
[…]
[inst2]organization=Department of Infomation Engineering, University of Padova,addressline=Via Gradenigo 6/B, city=Padova, postcode=35131, country=Italy
1 Introduction
@@
Autonomous agents operating in dynamic environments require robust training methods, with Reinforcement Learning (RL) offering a potent framework for such continuous adaptation. Control strategies are of paramount importance in domains characterized by continuous action spaces, as discussed in recent literature [1]. Actor-critic methods, particularly those utilizing temporal difference learning [2], represent a foundational approach within this area. However, the integration of Q-learning with deep neural networks to form a critic function has set new benchmarks [3], enhancing policy optimization through advanced policy gradient techniques [4, 5, 6].
@@
These state-of-the-art approaches, however, are not without their pitfalls, particularly the tendency of Q-learning to induce overestimation bias [7]. To combat this, the Twin Delayed Deep Deterministic Policy Gradient (TD3) method was developed, innovatively applying Clipped Double Q-Learning (CDQ) to refine accuracy in action value estimations [5]. By deploying two independently trained neural networks, this method systematically lowers the risk of overestimating the Q-values by adopting the lesser of the two during the training process.
Despite its benefits, CDQ can introduce underestimation bias, albeit with a lesser impact on policy modifications when compared to overestimation, as argued by the proponents of TD3 [5]. Furthermore, the introduced estimator leaves very little room for tuning the trade-off. Indeed, new developments [8] built on top of TD3 optimize for a convex combination of the minimum between the estimates, the CDQ estimate, and the maximum of the two.
@@
Recently, new algorithms further increased the pool of Q-networks in order to have a less noisy estimate of the true Q-learning target [9], getting to the point of using $10$ or more networks. These new enlarged pools of independent estimates are usually combined with multiple steps of updates, overshadowing previous algorithms’ performances. Instead, TD3 performs a single update per step. However, these enhancements introduce a very noticeable overhead, using up to $10$x more memory and $10$x more computational resources than the original algorithm they build on top of.
@@
The increase in computational cost of these new algorithms is prohibitive in many low-resource scenarios, for example, in edge computing applications. Many edge computing applications can benefit from data-driven-based approaches. For example, in the domain of autonomous driving, real-time sensing, and decision processing can be conducted by deploying edge computing nodes on self-driving vehicles; this advancement enables a reduction in response times and enhances driving safety [10]. Similarly, in the field of smart healthcare, edge computing nodes utilized on wearable and medical devices allow for the monitoring of patient’s physiological parameters in real-time. The collected data is then transmitted to the cloud, where it is analyzed and used for diagnosis, facilitating the realization of telemedicine and personalized medicine [11].
The ability to deploy deep learning models on the edge is lately gaining always more attention due to the appealing properties of on-device learning, for its decentralized computation, for economic purposes, and for its privacy preserving nature. Such a problem can be tackled using quantization [12] or pruning [13].
However, for the case of Reinforcement Learning, in order to allow the personalization of such models to the context in which the edge device is deployed, it is important to also tackle the algorithmic side. Indeed, differently from most other areas of machine learning, Deep RL requires additional computation and additional models for the learning to be carried out, thus adding a non-negligible overhead.
@@
In this paper, we propose an alternative to DDPG that tackles overestimation with a single $Q$ estimate. This algorithm, called Edge Delayed Deep Deterministic Policy Gradient (EdgeD3), has a computational cost lower than DDPG by $25\%$ while maintaining the same memory footprint. It does it with a new expectile loss, which induces an underestimation bias that evens the overestimation caused by Q-learning. Even though being cheaper, EdgeD3 has performances comparable or superior to the state-of-the-art in most cases, while using $30\%$ less processing resources and having $30\%$ less memory footprint compared to such more advanced methods. Thanks to its memory efficiency and requiring less computational resources, it is much more suited for low-resource settings, such as edge computing applications, where CPU computing time, energy savings, and memory usage are highly impactful. Furthermore, such an ability to require less computing and memory can allow for in-device learning, preserving privacy, which is ever more important in such scenarios. In addition, the introduced loss formulation allows for more control over the estimation bias with respect to the CDQ approach.
To benchmark and compare the proposed algorithm, we use a selection of Mujoco [14] robotics environments from the OpenAI Gym suite [15].
@@
The paper is structured as follows: in Section 2, we discuss recent contributions that are relevant to the scope of this paper, while in Section 3, we briefly review the required theoretical background. Section 4 discusses the newly introduced loss and its theoretical foundation. In Section 5 we discuss the stabilization of the optimization procedure for the DDPG algorithm and we introduce the final EdgeD3 algorithm. Section 6 compares the proposed EdgeD3 algorithm to the state-of-the-art both from a computational expensiveness side and performance point-of-view. Finally, Section 7 concludes the paper by briefly discussing some potential future lines of contributions.
2 Related work
@@
In the context of controlling continuous action spaces, recent contributions have tackled the dual issues of underestimation and overestimation biases by ensembling Q function estimates [19, 20, 21, 9]. Truncated Quantile Critics (TQC) was proposed by Kuznetsov et al. [19], which extends the Soft Actor-Critic (SAC) by integrating an ensemble of five critic estimates. This model not only offers a distributional representation of the critic but also implements truncation during critic updates to reduce overestimation bias. Moreover, TQC refines the actor update mechanism by employing an averaged ensemble of Qs.
@@
Remark 1.
The majority of contributions discussed in this section enhance TD3 or SAC by employing large ensembles of $Q$ Networks in the critic or by performing multiple steps of training at each time step, thereby introducing extra computational complexity. In contrast, our work concentrates on enhancing the performances by tackling the overestimation without adding additional computational burden on top of the original algorithm DDPG, thus being even cheaper than the original CDQ mechanism.
3 Background
@@
4 Overestimation and Underestimation in Q-Learning
@@
4.1 Tackling Overestimation with a single $Q$ estimate
@@
[…]
@@
[…]
@@
with $Z=\max(\alpha,\beta)$. Thanks to theorem 1, we can prove that $\alpha$ and $\beta$ control the overestimation-underestimation bias, namely, for a generic function $f$:
@@
In fig. 1, we show the function learned by a fourth-degree polynomial minimizing the expectile loss varying $\alpha,\beta$ approximating the function $y=0.1x^{3}+\epsilon,\,\,\epsilon\sim N(0,8)$. Features were sampled uniformly from $U(-10,10)$, and both features and targets were normalized in the range $\left[0,1\right]$. The parameters are initialized to 00 and optimized using Adam[25] with $0.001$ stepsize. Different $\alpha,\beta$ tradeoffs learn different expectiles of the distribution $p(y|x)$.
@@
As mentioned earlier, we add a normalizing constant $Z$ in front of the equation in order to have a fair comparison between algorithms. From an optimization standpoint, it’s equivalent to a change in the step size of the optimizer. Indeed, thanks to $Z$, the type of error we prefer to penalize, the one with the highest coefficient, has exactly $1$ as a constant in front, leading to an update that is equivalent to the original method. For the other type of error, on the other hand, the loss is multiplied by a constant $<1$, leading to a lower step size. This way, we can guarantee that the improvements shown by this proposal are due to the effectiveness of the loss and not by bigger step sizes induced by hidden constants in the loss.
Furthermore, we define a scheduled decay function $\lambda(t)$ with the property that $\lim_{t\rightarrow+\infty}\lambda(t)=1$. At every step, we will decay the distance between $\alpha,\beta$ in the following way:
@@
Figure 2: Training progress curves for continuous control tasks in OpenAI Gym, showing the effect of different choices of $\alpha,\beta$ in EdgeDDPG. Plots and shaded areas indicate mean and half a standard deviation, respectively from evaluation across 10 trials. Benchmarks were performed on 10 random seeds for simulator and network initializations. Curves are smoothed uniformly for visual clarity.
@@
A complete description of the algorithm can be found in algorithm 1. The new objective eq. 12 lays the ground for the final and proposed algorithm, EdgeD3, which will be discussed in Section 5.
@@
Figure 2 furthermore shows the tradeoff between overestimation and underestimation, highlighting how both of them can be detrimental. Indeed, it shows how preferring underestimation over overestimation has some diminishing returns eventually, so pushing for a strong underestimation bias can be as detrimental as accepting the overestimation bias of Q-learning. In Appendix appendix B, we report further tests of the algorithm on additional tasks and with different choices of $\alpha$ and $\beta$.
@@
Algorithm 1 Edge Deep Deterministic Policy Gradient (EdgeDDPG)
  Given $\alpha,\beta,\tau_{1},\tau_{2}$ and $\lambda(t)$
  Initialize critic $Q_{\theta}$, and actor $\mu_{\phi}$ networks
  Initialize replay memory $\mathcal{B}$
5:  repeat
     repeat
        $d=s^{\prime}\text{ is terminal}$
        Store $(s,a,r,s^{\prime},d)$ tuple in $\mathcal{B}$
10:        Sample mini-batch of $N$ tuples $(s,a,r,s^{\prime},d)$ from $\mathcal{B}$
        $y=r+\gamma Q_{\theta^{\prime}}(s^{\prime},\mu_{\phi^{\prime}}(s^{\prime}))\cdot(1-d)$
        Update $Q_{\theta}$ via GD using $\nabla L(\theta)$
        Update $\phi$ by deterministic policy gradient:
15:        $\nabla_{\phi}J(\phi)=\frac{1}{N}\Sigma\nabla_{a}Q_{\theta}(s,a)|_{a=\mu_{\phi}(s)}\nabla_{\phi}\mu_{\phi}(s)$
        Update target networks:
        $\theta^{\prime}\leftarrow\tau_{1}\theta+(1-\tau_{1})\theta^{\prime}$
        $\phi^{\prime}\leftarrow\tau_{2}\phi+(1-\tau_{2})\phi^{\prime}$
     until $d$ is false
20:     $\min(\alpha,\beta)\leftarrow\min(\alpha,\beta)+|\alpha-\beta|\cdot\lambda(t)$
  until $t<T$
5 Smoothing the optimization landscape

### v2

@@ Abstract
Deep Reinforcement Learning (DRL) has emerged as a powerful paradigm for learning complex policies directly from high-dimensional input spaces, enabling advances across a variety of domains. Modern DRL algorithms often rely on dual-network Q-learning architectures to approximate optimal policies to overcome overestimation bias. Recent research has introduced approaches leveraging multiple Q-functions to further mitigate overestimation effects and enhance policy reliability. However, there is a growing emphasis on deploying DRL in edge scenarios, where privacy concerns and stringent hardware constraints necessitate highly efficient algorithms. In such environments, the computational and memory efficiency of learning methods is of critical importance. In this context, we propose Edge Delayed Deep Deterministic Policy Gradient (EdgeD3), a novel reinforcement learning algorithm specifically designed for edge computing settings. EdgeD3 offers significant reductions in GPU time (by 25%) and computational and memory usage (by 30%), while consistently achieving or surpassing the performance of state-of-the-art algorithms across multiple benchmarks and in real-world tasks.
Note to Practitioners
Driven by the growing need for efficient computational solutions in automation, this research introduces the Edge Delayed Deep Deterministic Policy Gradient (EdgeD3), a novel Deep Reinforcement Learning algorithm designed for applications on edge devices with a limited computational budget. EdgeD3 is developed to enable on-device execution of policy learning, when computing resources are at a premium, such as in smart manufacturing and autonomous vehicle systems. EdgeD3 can learn highly effective policies, on par of state-of-the-art algorithms, while utilizing significantly fewer resources. This would empower autonomous devices to operate independently of cloud-based systems, fostering faster operational speeds and enhancing data privacy.
I Introduction
@@
Autonomous agents operating in dynamic environments require robust training methods, with Reinforcement Learning (RL) offering a potent framework for such continuous adaptation. Control strategies are of paramount importance in domains characterized by continuous action spaces, as discussed in recent literature [1].
Such approaches have been applied to nagivation [2, 3, 4], robotic manipulation [5, 6] and Athletic Intelligence [7]. Furthermore, they are also gaining attention in the multi-agent settings, controlling independently multiple agents [8, 9, 10].
Actor-critic methods, particularly those utilizing temporal difference learning [11], represent a foundational approach within this area. However, the integration of Q-learning with deep neural networks to form a critic function has set new benchmarks [12], enhancing policy optimization through advanced policy gradient techniques [13, 14, 15].
@@
These state-of-the-art approaches, however, are not without their pitfalls, particularly the tendency of Q-learning to induce overestimation bias [16]. To combat this, the Twin Delayed Deep Deterministic Policy Gradient (TD3) method was developed, innovatively applying Clipped Double Q-Learning (CDQ) to refine accuracy in action value estimations [14]. By deploying two independently trained neural networks, this method systematically lowers the risk of overestimating the Q-values by adopting the lesser of the two during the training process. Despite its benefits, CDQ can introduce underestimation bias, albeit with a lesser impact on policy modifications when compared to overestimation, as argued by the proponents of TD3 [14]. Furthermore, the introduced estimator leaves very little room for tuning the trade-off. Indeed, new developments [17] built on top of TD3 optimize for a convex combination of the minimum between the estimates, the CDQ estimate, and the maximum of the two. Thanks to such improvements, these algorithms have been widely applied to various sectors, such as robotic grasping [18, 19], wind turbine control [20], and path following [21].
@@
Recently, new algorithms further increased the pool of Q-networks in order to have a less noisy estimate of the true Q-learning target [22, 23, 24, 25, 26], using up $10$ or more networks. These new enlarged pools of independent estimates are usually combined with multiple steps of updates, overshadowing previous algorithms’ performances, while TD3[14] performs a single update per step. These enhancements introduce a very noticeable overhead, using up to 1010x more memory and 1010x more computational resources than the original algorithm they build on top of.
@@
The increase in computational cost of these new algorithms [22, 23, 24, 25, 26] is prohibitive in many low-resource scenarios, for example, in edge computing applications. Many edge computing applications can benefit from data-driven approaches. For example, in the domain of autonomous driving, real-time sensing and decision processing can be conducted by deploying edge computing nodes on self-driving vehicles; this advancement enables a reduction in response times and enhances driving safety [27]. Similarly, in the field of smart healthcare, edge computing nodes utilized on wearable and medical devices allow for the monitoring of patients’ physiological parameters in real-time. The collected data is then transmitted to the cloud, where it is analyzed and used for diagnosis, facilitating the realization of telemedicine and personalized medicine [28]. The ability to deploy deep learning models on the edge is lately gaining more and more attention due to the appealing properties of on-device learning, for its decentralized computation, for economic purposes, and for its privacy-preserving nature. Such a problem can be tackled using quantization [29] or pruning [30]. However, for the case of Reinforcement Learning, in order to allow the personalization of such models to the context in which the edge device is deployed, it is also important to tackle the algorithmic side. Indeed, differently from most other areas of machine learning, Deep RL requires additional computation and additional models for the learning to be carried out, thus adding a non-negligible overhead. Many applications [2, 3, 4, 5, 6, 8, 9, 10] can be found using these approaches to learn effective policies, and notably most of them share the property that a real world deployment of the proposed approach would imply a constrained computational reosurce availability. For this reason, it is of paramount importance to consider the computational cost of Deep RL algorithms, even over raw policy-performance, which is an aspect drastically underrated by these latest advancements in continuous control.
@@
In this paper, we propose an alternative to Deep Deterministic Policy Gradient (DDPG) that tackles overestimation with a single $Q$ estimate. This algorithm, called Edge Delayed Deep Deterministic Policy Gradient (EdgeD3), has a computational cost lower than DDPG by $25\%$ while maintaining the same memory footprint. It does so with a new expectile loss, which induces an underestimation bias that evens the overestimation caused by Q-learning. Even though being cheaper, EdgeD3 has performances comparable or superior to the state-of-the-art in most cases, while using $30\%$ less processing resources and having $30\%$ less memory footprint compared to such more advanced methods. Thanks to its memory efficiency and requiring less computational resources, it is much more suited for low-resource settings, such as edge computing applications, where CPU computing time, energy savings, and memory usage are highly impactful. Furthermore, such an ability to require less computing and memory can allow for in-device learning, preserving privacy, which is ever more important in such scenarios. In addition, the introduced loss formulation allows for more control over the estimation bias with respect to the CDQ approach. To benchmark and compare the proposed algorithm in simulation, we use a selection of Mujoco [31] robotics environments from the OpenAI Gym suite [32]. Furthermore, we also validate the proposed approach on real-world navigation tasks, using custom TurtleBots, which are widely used mobile robots [33, 34, 35], with limited computing resources.
The key contributions of this paper are summarized as follows:
  • We propose a novel approach to tackle overestimation in DDPG-like algorithms using Expectile estimations, thus not requiring additional components;
  • We show that such a quantity can be efficiently estimated via samples, thus allowing convergence under simulation-based learning;
  • We empirically evaluate the proposed approach on different robotic simulation tasks to estimate the improvement brought by the novel loss formulation;
  • We test the proposed approach on two different real-world navigation tasks to highlight the importance of update frequency.
@@
The paper is structured as follows: in Section II, we discuss recent contributions that are relevant to the scope of this paper, while in Section III, we briefly review the required theoretical background. Section IV discusses the newly introduced loss and its theoretical foundation. In Section V we discuss the stabilization of the optimization procedure for the DDPG algorithm and we introduce the final EdgeD3 algorithm. Section VI compares the proposed EdgeD3 algorithm to the state-of-the-art both from a computational expensiveness side and performance point-of-view, both on simulation and on two real-world navigation tasks. Finally, Section VII concludes the paper by briefly discussing some potential future lines of contributions.
II Related work
@@
In the context of controlling continuous action spaces, recent contributions have tackled the dual issues of underestimation and overestimation biases by ensembling Q function estimates [39, 23, 40, 22, 24, 25, 26]. Truncated Quantile Critics (TQC) was proposed by Kuznetsov et al. [39], which extends the Soft Actor-Critic (SAC) by integrating an ensemble of five critic estimates. This model not only offers a distributional representation of the critic but also implements truncation during critic updates to reduce overestimation bias. Moreover, TQC refines the actor update mechanism by employing an averaged ensemble of Qs.
@@
Recent advancements [24, 25, 26] have introduced innovative methods for aggregating Q-value predictions to achieve more accurate and unbiased, non-overestimated evaluations of action quality. In contrast, [25] proposes a distinct methodology that leverages an energy-based model (EBM) to address the overestimation problem. However, this approach necessitates the use of sampling procedures to obtain the final estimate, rendering it impractical for deployment in resource-constrained environments.
Remark: the contributions discussed in this section enhance TD3 or SAC by employing large ensembles of $Q$ Networks in the critic or by performing multiple steps of training at each time step, thereby making them infeasible for edge scenarios. In contrast, our work concentrates on enhancing the performance by tackling overestimation bias, further improving over the CDQ mechanism, while reducing computational burden of the DDPG algorithm.
III Background
@@
IV Overestimation and Underestimation in Q-Learning
@@
IV-A Tackling Overestimation with a single $Q$ estimate
@@
$\begin{split}\frac{\alpha}{2\max(\alpha,\beta)}\int_{-\infty}^{t}&(t-x)dF(x)=\\ &\frac{\beta}{2\max(\alpha,\beta)}\int_{t}^{+\infty}(x-t)dF(x)\end{split}$ (10)
@@
Dividing both sides by $(1-\tau)$ and then by $\int_{t}^{+\infty}(x-t)dF(x)$, we can rewrite eq. 9 in the following way:
$\frac{\tau}{1-\tau}=\frac{\int_{-\infty}^{t}(t-x)dF(x)}{\int_{t}^{+\infty}(x-t)dF(x)}$
@@
with $Z=\max(\alpha,\beta)$, $y$ being the target and $f_{\theta}(x)$ being the regressor network. Thanks to theorem 1, we can prove that $\alpha$ and $\beta$ control the overestimation-underestimation bias, namely, denoting by $f^{\alpha,\beta}(x)$ the optimal solution to $L^{\alpha,\beta}(f(x),y)$:
@@
In fig. 1, we show the function learned by a fourth-degree polynomial minimizing the expectile loss varying $\alpha,\beta$ approximating the function $y=0.1x^{3}+\epsilon,\,\,\epsilon\sim N(0,8)$. Features were sampled uniformly from U(10,10)U(-10,10), and both features and targets were normalized in the range $\left[0,1\right]$. The parameters are initialized to 0 and optimized using Adam[44] with $0.001$ stepsize. Different $\alpha,\beta$ tradeoffs learn different expectiles of the distribution $p(y|x)$.
@@
As mentioned earlier, we add a normalizing constant $Z$ in front of the equation in order to have a fair comparison between algorithms. From an optimization standpoint, it’s equivalent to a change in the step size of the optimizer. Indeed, thanks to $Z$, the type of error we prefer to penalize, the one with the highest coefficient, has exactly $1$ as a constant in front, leading to an update that is equivalent to the original method. For the other type of error, on the other hand, the loss is multiplied by a constant <1<1, leading to a lower step size. This way, we can guarantee that the improvements shown by this proposal are due to the effectiveness of the loss and not by bigger step sizes induced by hidden constants in the loss.
Corollary 1.1 (Asymptotic convergence of Q-learning with Expectile loss).
Given a scheduled decay function $\lambda(t)$, such that $\lim_{t\rightarrow+\infty}\lambda(t)=1$, we can directly reduce the induced bias by the expectile loss, by applying at each step:
@@
Remark 1 (Consistency and Asymptotic Normality of Empirical Expectiles).
Let $\{X_{i}\}_{i=1}^{n}$ be i.i.d. samples from a distribution with cumulative distribution function $F$, and let $t_{n}$ be the empirical $\tau$-expectile obtained by minimizing the sample counterpart of Equation 9. Then, under standard regularity conditions for M-estimators, $t_{n}$ is a consistent estimator of the true $\tau$-expectile $t^{*}$ (i.e., $t_{n}\xrightarrow{p}t^{*}$), and it satisfies a central limit theorem:
$\sqrt{n}(t_{n}-t^{*})\xrightarrow{d}\mathcal{N}\left(0,\;\left[\Psi(t^{*},\tau)\right]^{-2}\right),$
where $\Psi(t,\tau)$ is the derivative of the population objective function (eq. 9) associated with the expectile definition at $t$.
Corollary 1.2 (Monte Carlo Estimation of Expectile Function).
When the integrals in Equation 9 are approximated via Monte Carlo sampling,
$\hat{t}=\arg\min_{t}\sum_{i=1}^{n}\omega_{i}\cdot\ell_{\tau}(t,X_{i}),$
where $\ell_{\tau}(\cdot,\cdot)$ represents the asymmetric loss defining the expectile, and $\omega_{i}$ are weights (often uniform, $\omega_{i}=\frac{1}{n}$), then $\hat{t}$ inherits the consistency and asymptotic normality as described in Theorem 1. Consequently,
$\sqrt{n}(\hat{t}-t^{*})\xrightarrow{d}\mathcal{N}\left(0,\;\left[\Psi(t^{*},\tau)\right]^{-2}\right).$
For more information and formal proofs of remark 1 and corollary 1.2 refer to [43, 46].
Remark 2.
These theoretical results validate the use of the expectile loss in place of traditional MSE-based updates by confirming that the estimator converges to the correct level of under- or over-estimation specified by the parameters $\alpha$ and $\beta$ (or equivalently $\tau$). These properties also justify the robustness and adaptability of expectile-based methods in function approximation and value estimation within reinforcement learning frameworks.
To conclude, theorem 1 shows how the proposed loss can directly manipulate the wanted overestimation/underestimation, while corollary 1.1 show how to prove the convergence of the proposed method, ensuring no bias will be introduced at convergence. Finally, remark 1 and corollary 1.2 show that the expectile loss eq. 11 allows a consistent and asymptotic convergence to the true expectile, allowing theorem 1 to hold also in a setting where Monte Carlo estimation is employed.
Figure 2: Training progress curves for continuous control tasks in OpenAI Gym, showing the effect of different choices of $\alpha,\beta$ in EdgeDDPG. Plots and shaded areas indicate mean and standard deviation, respectively from evaluation across 10 trials. Benchmarks were performed on 10 random seeds for simulator and network initializations. Curves are smoothed uniformly for visual clarity.
@@
A complete description of the algorithm can be found in section -A. The new objective eq. 12 lays the ground for the final and proposed algorithm, EdgeD3, which will be discussed in Section V.
@@
Figure 2 furthermore shows the tradeoff between overestimation and underestimation, highlighting how both of them can be detrimental. Indeed, it shows how preferring underestimation over overestimation has some diminishing returns eventually, so pushing for a strong underestimation bias can be as detrimental as accepting the overestimation bias of Q-learning. In appendix B, we report further tests of the algorithm on additional tasks and with different choices of $\alpha$ and $\beta$.
@@
V Smoothing the optimization landscape